### v1

# Fabian Mauthe

Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Hagmeyer, S.; Mauthe, F.; Zeiler, P. (2021): Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087.
Fabian Mauthe
Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
Goeppingen, Germany
fabian.mauthe@hs-esslingen.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Stuttgart, Germany
christopher.braun@iff.uni-stuttgart.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Stuttgart, Germany
julian.raible@iff.uni-stuttgart.de
[…]
Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
Goeppingen, Germany
peter.zeiler@hs-esslingen.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Fraunhofer Institute of Manufacturing Engineering and Automation IPA
Stuttgart, Germany
marco.huber@ieee.org
1 Prognostics and Health Management
In the realm of modern manufacturing, significant emphasis is placed on improving the reliability, performance, and service life of complex engineering systems, which has catalyzed the emergence of Prognostics and Health Management (PHM) as an essential discipline in the past two decades [1, 2]. PHM includes a range of methods dedicated to the early detection, prognosis, and mitigation of impending faults or failures in ESs, thereby facilitating sophisticated maintenance strategies and enhancing operational efficiency [3].
Central to the efficacy of PHM methods is the acquisition and analysis of degradation data, which encapsulates the evolving health condition of ESs over time [3]. Degradation data serves as a rich source of information, offering invaluable insights into the underlying degradation processes, failure modes, and performance trends of ESs [2, 4]. Through the integration of advanced sensing technologies, data analytics, and machine learning algorithms, researchers and practitioners can harness degradation data to develop prognostic models, optimize maintenance schedules, and make informed operational decisions [1].
In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview of publicly available degradation data sets is presented as well as briefly described. The authors aim to update the overview regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
2 Tasks of Diagnostics and Prognostics
Extensive and representative degradation data enables addressing a specific task—prognosis—within PHM, with the subdivision of PHM into distinct tasks. Based on [1], [5], and [4], the tasks of diagnostics and prognostics within PHM can be subdivided as follows:
  • Fault detection/anomaly detection: Detect a fault state or anomaly in an ES without considering its root cause. This results in a binary classification problem with the states fault or no fault.
  • Diagnosis: Assign one or more causes to a detected fault state.
  • Health assessment: Assess the state of degradation or the current risk of failure of an ES based on its current condition.
  • Prognosis: Prediction of the future degradation behavior or the current Remaining Useful Life (RUL).
The prognosis task poses the greatest challenge in diagnostics and prognostics within PHM [6]. This difficulty stems from the intricate nature of failure mechanisms and the multitude of factors influencing degradation progression. Therefore, sufficient degradation data is of huge importance for this task and needs to be comprehensively available.
3 Data Sets for Diagnostics and Prognostics
Table Appendix: Overview of Publicly Available Data Sets in the Appendix: Overview of Publicly Available Data Sets contains an overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically. This is the extended overview from [4], which already contains 70 data sets.
The designation of the data sets is divided into two parts, consisting of the source or platform and the title under which they were published. In addition, the overview contains the following columns:
  • Domain and Application: The associated domain from which the data set originates and the respective application within this domain.
  • Task: The corresponding task within PHM addressable with this data set.
  • URL: The respective hyperlink to the source or platform hosting the data set.
4 Description of Data Sets
The overview of publicly available degradation data sets in Table Appendix: Overview of Publicly Available Data Sets comprises a total of 79 data sets. These data sets originate from eleven different domains. The domains as well as the applications presented do not claim to be exhaustive and are based solely on the available data sets. With 17 assigned data sets, the Mechanical component domain is the most extensive, followed by the Electrical component domain with 13 assigned data sets. The two most frequently occurring applications, Bearing (ten) and Battery (six), also originate from these two domains. In total, the domains contain 36 different applications.
Except for the Building domain, which is sparsely represented in PHM, the data sets are evenly distributed for the remaining domains, with at least four and at most nine assignments per domain. With only one data set for each application, the Mechatronic system domain contains the greatest variation of applications. Moreover, there are five data sets with unknown applications and therefore, no domain can be assigned. Data sets with degradation data resulting from a simulated application are assigned to the corresponding domains but marked. With 32 assigned data sets for prognosis and 29 for diagnosis, these tasks are the most represented. For fault detection, 15 data sets are assigned, and for health assessment, 3 data sets are assigned. The Electrical component domain contains the most data sets (ten) for the task prognosis, followed by the Drive technology domain with six and the Manufacturing process as well as Process technology domains with four assigned data sets each.
Of the 17 assigned data sets in the Mechanical component domain, only two are for task prognosis. However, for the task diagnosis, the Mechanical component domain contains the most data sets with eleven assignments. Followed by the Mechatronic system domain with five and the Robotic domain with four assigned data sets. Regarding fault detection, the domains Mechanical component and Production system contain the most data sets with four assignments each. With two data sets with an Unknown domain and one data set from the Drive technology domain, the overview contains the fewest data sets for the health assessment task.
Appendix: Overview of Publicly Available Data Sets
Overview of publicly available data sets
No.
Data Set Designation
Reference to Data Origin
Domain
Application
Task URL
[…]
Table 0 – continued from previous page
No.
Data Set Designation
Reference to Data Origin
Domain
Application
Task URL
[…]
Continued on next page
[…]
1
AIDAR Lab - Air Compressor
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2016). Intelligent Condition Based Monitoring using Acoustic Signals for Air Compressors, IEEE Transactions on Reliability, vol. 65, no. 1, pp. 291-309.
Mechatronic system
Air Compressor
Diagnosis
Link
2
AIDAR Lab - Drill Bit
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26.
Manufacturing process
Drill bit
Diagnosis
Link
3
Aramis - Data Challenge ESREL2020PSAM15
Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France).
Unknown
Simulation (time-continuous stochastic process)*
Fault detection
Link
4
Backblaze - Hard Drive Stats
Backblaze, Inc.
Drive technology
Disk Drives
Fault detection
Link
5
Bearing Data Center - Fan and Bearing
Case Western Reserve University, USA
Mechanical component
Bearing
Diagnosis
Link
6
GitHub - Predictive Maintenance using PySpark
GitHub User: linya9191
Simulation
Simulation*
Fault detection
Link
7
Kaggle - Air pressure system failures in Scania trucks
Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm
Mechatronic system
Air Pressure System
Fault detection
Link
8
Kaggle - CNC Mill Tool Wear
System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan
Manufacturing process
Mill Tool
Fault detection
Link
9
Kaggle - Condition Data with Random Recording Time
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Prognosis
Link
10
Kaggle - E-coating ultrafiltration maintenance
Tinsley Bridge Limited, Sheffield, United Kingdom
Manufacturing process
Electrophoresis Painting
Prognosis
Link
11
Kaggle - Genesis Demonstrator
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Robotic
Pick-and-Place Demonstrator
Diagnosis
Link
12
Kaggle - Microsoft Azure Predictive Maintenance
Azure AI Notebooks for Predictive Maintenance
Unknown
Unknown Machine
Fault detection
Link
13
Kaggle - One Year Industrial Component Degradation
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
Shrink-Wrapper
Fault detection
Link
14
Kaggle - Kaggle - Predictive Maintenance 1
Creators Yuan Yao and Zhao Yuqi
Production system
Log Data
Fault detection
Link
15
Kaggle - Preventive to Predictive Maintenance
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Prognosis
Link
16
Kaggle - Production Plant Data for Condition Monitoring
von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018). Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems. Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRP-CMS
Production system
Productionline
Fault detection
Link
17
Kaggle - Prognosis based on Varying Data Quality
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Prognosis
Link
18
Kaggle - Pump
Kaggle User: UnknownClass
Process technology
Water Pump
Fault detection
Link
19
Kaggle - Sensor Fault Detection
Schneider-Electric, France
Electrical component
Temperature Sensor
Fault detection
Link
20
Kaggle - Versatile Production System
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
Versatile Production System
Fault detection
Link
21
Mendeley - Brushless DC motor
Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019). Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2
Drive technology
Brushless DC Motor
Prognosis
Link
22
Mendeley - Diesel Engine Faults
Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil
Drive technology
Diesel Engine
Diagnosis
Link
23
Mendeley - HUST Bearing
Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3
Mechanical component
Bearing
Diagnosis
Link
24
MFPT - Condition Based Maintenance Fault
Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems
Mechanical component
Bearing
Diagnosis
Link
25
NASA - Accelerated Battery Life Testing
Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F. (2023). Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks. Proccedings of the Annual Conference of the PHM Society, 15(1). https://doi.org/10.36001/phmconf.2023.v15i1.3463
Electrical component
Battery
Prognosis
Link
26
NASA - Bearing Data Set
J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. "Bearing Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Mechanical component
Bearing
Prognosis
Link
27
NASA - Capacitor Electrical Stress
J. Renwick, C. Kulkarni, and J Celaya "Capacitor Electrical Stress Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Electrolytic Capacitors
Prognosis
Link
28
NASA - Capacitor Electrical Stress 2
J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel "Capacitor Electrical Stress Data Set - 2", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Electrolytic Capacitors
Prognosis
Link
29
NASA - CFRP Composites Data Set
Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang "CFRP Composites Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Materials
Carbon Fibre-Reinforced Polymer Composite
Prognosis
Link
30
NASA - HIRF Battery
C. Kulkarni, E. Hogge, C. Quach and K. Goebel "HIRF Battery Data Set", NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognostic-data-repository), NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Prognosis
Link
31
NASA - IGBT
J. Celaya, Phil Wysocki, and K. Goebel "IGBT Accelerated Aging Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Transistor
Prognosis
Link
32
NASA - Li-ion Battery Aging Datasets
B. Saha and K. Goebel. "Battery Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Prognosis
Link
33
NASA - Milling Data set
A. Agogino and K. Goebel. BEST lab, UC Berkeley. "Milling Data Set ", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Manufacturing process
Milling
Progrognosis
Link
34
NASA - MOSFET Thermal Overstress Aging
J. R. Celaya, A. Saxena, S. Saha, and K. Goebel "MOSFET Thermal Overstress Aging Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Transistor
Prognosis
Link
35
NASA - Randomized Battery Usage Data Set
B. Bole, C. Kulkarni, and M. Daigle "Randomized Battery Usage Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Prognosis
Link
36
NASA - Small Satellite Power Simulation
C. Kulkarni and A. Guarneros "Small Satellite Power Simulation Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Diagnosis
Link
37
NASA - Turbofan engine degradation simulation data set
A. Saxena and K. Goebel. "Turbofan Engine Degradation Simulation Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
Aircraft Engine*
Prognosis
Link
38
NASA - Turbofan engine degradation simulation data set 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). "Aircraft Engine Run-to-Failure Dataset under real flight conditions", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
Aircraft Engine*
Prognosis
Link
39
NIST - Robot Arm Position Accuracy
National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division
Robotic
Industrial Robot
Diagnosis
Link
40
OpenEI - Evaluation of Building Fault
Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019)
Building
Building
Diagnosis
Link
41
OpenEI - Gearbox Fault Diagnosis
Yogesh Pandy, National Renewable Energy Laboratory
Mechanical component
Gear
Fault detection
Link
42
PHM Data Challenge 2008 - Turbofan
NASA Prognostics Center of Excellence
Drive technology
Aircraft Engine*
Prognosis
Link
43
PHM Data Challenge 2009 - Gearbox Fault Detection
PHM Society, Gearbox fault detection data set, 2010
Mechanical component
Gear System (Gears, Bearings, Shafts)
Diagnosis
Link
44
PHM Data Challenge 2010 - CNC milling machine cutters
X. Li, B.S. Lim, J.H. Zhou, S. Huang, S.J. Phua, K.C. Shaw, and M.J. Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798
Manufacturing process
CNC Milling Machine Cutters
Prognosis
Link
45
PHM Data Challenge 2011 - Anemometer Fault Detection
Creators Unknown
Mechanical component
Anemometer
Fault detection
Link
46
PHM Data Challenge 2013 - Unknown
Creators Unknown
Unknown
Unknown
Health assessment
Link
47
PHM Data Challenge 2014 - Unknown
Creators Unknown
Unknown
Unknown
Health assessment
Link
48
PHM Data Challenge 2015 - Plant Fault Detection
Creators Unknown
Production system
Industrial Plant Monitoring
Diagnosis
Link
49
PHM Data Challenge 2016 - Semiconductor CMP
Crystec Technology Trading GmbH
Manufacturing process
Chemicalmechanical Planarization System
Prognosis
Link
50
PHM Data Challenge 2017 - Bogie Vehicle
Creators Unknown
Mechanical component
Bogie
Diagnosis
Link
51
PHM Data Challenge 2018 - Ion Mill in Waver Manufacturing
Kai Goebel (at this time at NASA)
Production system
Ion Mill Etching Tool
Prognosis
Link
52
PHM Data Challenge 2019 - Fatigue Cracks
Creators Unknown
Materials
Aluminum Structure with Dynamic Loading
Prognosis
Link
53
PHM Data Challenge 2020 Europe - Filtration System
Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412.
Process technology
Filtration
Prognosis
Link
54
PHM Data Challenge 2021 - Turbofan 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). "Aircraft Engine Run-to-Failure Dataset under real flight conditions", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2)
Drive technology
Aircraft Engine*
Prognosis
Link
55
PHM Data Challenge 2021 Europe: SCARA-robot
Swiss Centre for Electronics and Microtechnology (CSEM)
Robotic
Industrial Robot
Diagnosis
Link
56
PHM Data Challenge 2022 - Rock Drills
Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden
Mechatronic system
Rock Drills
Diagnosis
Link
57
PHM Data Challenge 2022 Europe - PCB Production line
Danilo Giordano and Martino Trevisan, (2022). "PHME Data Challenge". European conference of the prognostics and health management society. / BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en
Production system
PCB Production Line
Diagnosis
Link
58
PHM Data Challenge 2023 - Gearbox
PHM Society, Gearbox data set, 2023
Mechanical component
Gear
Diagnosis
Link
59
PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
Japan Aerospace Exploration Agency (JAXA)
Drive technology
Propulsion System*
Health assessment
Link
60
PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
FEMTO-ST Institute, Besançon, France
Mechanical component
Bearing
Prognosis
Link
61
PHM IEEE Data Challenge 2014 - Fuel Cell
FCLAB Research Federation (FR CNRS 3539, France)
Electrical component
Fuel Cell
Prognosis
Link
62
PHM IEEE Data Challenge 2023 - Planetary Gearbox
Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC)
Mechanical component
Planetary Gear
Diagnosis
Link
63
PHM Society - Electromechanical Ball Screw Drive
GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy, under Award Number DE-AR0001290
Mechatronic system
Electromechanical Ball Screw Drive*
Diagnosis
Link
64
ResearchGate - Driveline Unbalanced Shaft
Giacomo Barbieri and David Sanchez-Londono and Laura Cattaneo and Luca Fumagalli and David Romero
Mechanical component
Unbalanced shaft (driveline)
Diagnosis
Link
65
SDOL - Diagnostics 101 bearing data
Kim, S., An, D., Choi, J-H. (2020). Diagnostics 101: A Tutorial for Fault Diagnostics of Rolling Element Bearing Using Envelope Analysis in MATLAB. Applied Sciences. 10(20):7302. doi:10.3390/app10207302
Mechanical component
Bearing
Fault detection
Link
66
SDOL - HS Gear
Creators Unknown
Mechanical component
Bearing
Fault detection
Link
67
SDOL - KAU Gear
Creators Unknown
Mechanical component
Bearing
Diagnosis
Link
68
Toyota Research Institute - Battery Cycle Life
Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-019-0356-8
Electrical component
Battery
Prognosis
Link
69
UCI - AI4I 2020 Predictive Maintenance
Matzka, S. (2020). “Explainable Artificial Intelligence for Predictive Maintenance Applications," Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp. 69-74, doi: 10.1109/AI4I49448.2020.00023 .
Production system
Manufacturing Machine
Diagnosis
Link
70
UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University
Drive technology
Naval Propulsion Plant*
Prognosis
Link
71
UCI - Condition monitoring of hydraulic systems Data Set
ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany
Mechatronic system
Hydraulic System
Prognosis
Link
72
UCI - Mechanical Analysis
F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy
Mechatronic system
Electromechanical Devices
Diagnosis
Link
73
UCI - Robot Execution Failures Data Set
Universidade Nova de Lisboa, Monte da Caparica, Portugal
Robotic
Industrial Robot
Diagnosis
Link
74
UCI - Steel Plates Faults
Semeion Research Center of Sciences of Communication - Rome, Italy
Materials
Steel Plates Faults
Diagnosis
Link
75
Uni Lulea - Vibration Signals Wind Turbines
Luleå University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements)
Mechanical component
Bearing (Wind Turbines)
Fault detection
Link
76
Uni Paderborn KAt - Bearing Damage
University Paderborn Kat
Mechanical component
Bearing
Diagnosis
Link
77
Virkler
Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153.
Materials
Aluminum Plate
Prognosis
Link
78
WASEDA - Fault Detection and Classification
De Bruijn, B., Nguyen, T. A., Bucur, D., & Tei, K. (2016). Benchmark datasets for fault detection and classification in sensor data. In A. Ahrens, O. Postolache, & C. Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Confererence on Sensor Networks, pp. 185-195. . SciTePress. https://doi.org/10.5220/0005637901850195
Electrical component
Sensors
Diagnosis
Link
79
Zenodo - Predictive maintenance dataset
Huawei Munich Research Center
Mechatronic system
Elevator
Prognosis
Link