### v2

# Fabian Mauthe

Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Fabian Mauthe, Luca Steinmann, and Peter Zeiler. Overview and analysis of publicly available degradation data sets for tasks within prognostics and health management. In Eirik Bjorheim Abrahamsen, Terje Aven, Frederic Bouder, Roger Flage, and Marja Yloenen, editors, 35th European Safety and Reliability Conference. (Submitted). Research Publishing, 2025.
Fabian Mauthe
Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
Goeppingen, Germany
fabian.mauthe@hs-esslingen.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Stuttgart, Germany
christopher.braun@iff.uni-stuttgart.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Stuttgart, Germany
julian.raible@iff.uni-stuttgart.de
[…]
Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
Goeppingen, Germany
peter.zeiler@hs-esslingen.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Fraunhofer Institute of Manufacturing Engineering and Automation IPA
Stuttgart, Germany
marco.huber@ieee.org
1 Prognostics and Health Management
In the realm of modern manufacturing, significant emphasis is placed on improving the reliability, performance, and service life of complex engineering systems, which has catalyzed the emergence of Prognostics and Health Management (PHM) as an essential discipline in the past two decades [1, 2]. PHM includes a range of methods dedicated to the early detection, prognosis, and mitigation of impending faults or failures in ESs, thereby facilitating sophisticated maintenance strategies and enhancing operational efficiency [3].
Central to the efficacy of PHM methods is the acquisition and analysis of degradation data, which encapsulates the evolving health condition of ESs over time [3]. Degradation data serves as a rich source of information, offering invaluable insights into the underlying degradation processes, failure modes, and performance trends of ESs [2, 4]. Through the integration of advanced sensing technologies, data analytics, and machine learning algorithms, researchers and practitioners can harness degradation data to develop prognostic models, optimize maintenance schedules, and make informed operational decisions [1].
In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview from [5] of publicly available degradation data sets is presented. The authors aim to update the overview regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
2 Tasks of Diagnostics and Prognostics
Extensive and representative degradation data enables addressing a specific task—prognosis—within PHM, with the subdivision of PHM into distinct tasks. Based on [1], [6], and [4], the tasks of diagnostics and prognostics within PHM can be subdivided as follows:
  • Fault detection/anomaly detection: Detect a fault state or anomaly in an ES without considering its root cause. This results in a binary classification problem with the states fault or no fault.
  • Diagnosis: Assign one or more causes to a detected fault state.
  • Health assessment: Assess the state of degradation or the current risk of failure of an ES based on its current condition.
  • Prognosis: Prediction of the future degradation behavior or the current Remaining Useful Life (RUL).
3 Data Sets for Diagnostics and Prognostics
Table Appendix: Overview of Publicly Available Data Sets in the Appendix: Overview of Publicly Available Data Sets contains an overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically. This is the extended overview from [4], presented in [5].
The designation of the data sets is divided into two parts, consisting of the source or platform and the title under which they were published. In addition, the overview contains the following columns:
  • Domain, Application, and Signals: The associated domain from which the data set originates and the respective application within this domain. In addition, the signals present in the respective data set are listed.
  • Task: The corresponding task within PHM addressable with this data set.
  • URL: The respective hyperlink to the source or platform hosting the data set.
The detailed PHM-specific analysis of the data sets can be found in [5].
Appendix: Overview of Publicly Available Data Sets
{xltabular}
|p0.6cm|X|p4.1cm|p2.3cm|p1.0cm| Overview of publicly available data sets
No. Data Set Designation
Reference to Data Origin Domain
Application and Signal Task URL
[…]
Table 0 – continued from previous page
No. Data Set Designation
Reference to Data Origin Domain
Application and Signal Task URL
[…]
  Continued on next page
[…]
1 4TU - Lifecycle ageing tests on commercial 18650 Li ion cell
Trad, Khiem (2021): Lifecycle ageing tests on commercial 18650 Li ion cell @ 25°C and 45°C. Version 1. 4TU.ResearchData. dataset. https://doi.org/10.4121/13739296.v1 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
2 4TU - Motor Current and Vibration Monitoring Dataset
Bruinsma, Sietze; Geertsma, Rinze; Loendersloot, Richard; Tinga, Tiedo (2024): Motor Current and Vibration Monitoring Dataset for various Faults in an E-motor-driven Centrifugal Pump. doi: 10.4121/2b61183ec14f-4131-829b-cc4822c369d0.v3 Process technology
Centrifugal pumps
Signals: current, vibration, voltage Diagnosis Link
3 AIDAR Lab - Air Compressor
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2016). Intelligent Condition Based Monitoring using Acoustic Signals for Air Compressors, IEEE Transactions on Reliability, vol. 65, no. 1, pp. 291-309. Mechatronic system
Air Compressor
Signals: acoustic emission Diagnosis Link
4 AIDAR Lab - Drill Bit
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26. Manufacturing process
Drill bit
Signals: vibration Diagnosis Link
5 Aramis - Data Challenge ESREL2020PSAM15
Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France). Unknown
Simulation (time-continuous stochastic process)*
Signals: vibration Fault detection Link
6 Backblaze - Hard Drive Stats
Backblaze, Inc. Drive technology
Disk Drives
Signals: temperature Fault detection Link
7 Bearing Data Center - Fan and Bearing
Case Western Reserve University, USA Mechanical component
Bearing
Signals: acoustic emission Diagnosis Link
8 Calce - Battery Data Repository
CALCE Center for Advanced Life Cycle Engineering, University of Maryland, website: https://calce.umd.edu/battery-data Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
9 GitHub - Predictive Maintenance using PySpark
GitHub User: linya9191 Simulation
Simulation*
Signals: unknown Fault detection Link
10 GitHub - XJTU-SY Bearing Datasets
Biao Wang, Yaguo Lei, Naipeng Li, Ningbo Li (2020). A Hybrid Prognostics Approach for Estimating Remaining Useful Life of Rolling Element Bearings, IEEE Transactions on Reliability, vol. 69, no. 1, pp. 401-412. DOI: 10.1109/TR.2018.2882682. Mechanical component
Bearing
Signals: vibration Prognosis Link
11 Kaggle - Air pressure system failures in Scania trucks
Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm Mechatronic system
Air Pressure System
Signals: anonymized Fault detection Link
12 Kaggle - Bearings with Varying Degradation Behaviors
Mauthe, F.; Hagmeyer, S.; Zeiler, P. (2025). Holistic simulation model of the temporal degradation of rolling bearings. In E. B. Abrahamsen, T. Aven, F. Bouder, R. Flage, and M. Yloenen (Eds.), Proceedings of the 35nd European Safety and Reliability Conference (Submitted). Research Publishing. Mechanical component
Bearing (simulated)
Signals: vibration Prognosis Link
13 Kaggle - CNC Mill Tool Wear
System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan Manufacturing process
Mill Tool
Signals: acceleration, current, feedrate, position, speed, voltage Fault detection Link
14 Kaggle - Condition Data with Random Recording Time
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
Filtration
Signals: flow rate, pressure Prognosis Link
15 Kaggle - E-coating ultrafiltration maintenance
Tinsley Bridge Limited, Sheffield, United Kingdom Manufacturing process
Electrophoresis Painting
Signals: flow rate, pressure, temperature Prognosis Link
16 Kaggle - Genesis Demonstrator
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Robotic
Pick-and-Place Demonstrator
Signals: acceleration, current, force, speed Diagnosis Link
17 Kaggle - Microsoft Azure Predictive Maintenance
Azure AI Notebooks for Predictive Maintenance Unknown
Unknown Machine
Signals: pressure, speed, vibration, voltage Prognosis Link
18 Kaggle - One Year Industrial Component Degradation
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
Shrink-Wrapper
Signals: position, speed, torque Fault detection Link
19 Kaggle - Predictive Maintenance 1
Creators Yuan Yao and Zhao Yuqi Production system
Log Data
Signals: unknown Fault detection Link
20 Kaggle - Preventive to Predictive Maintenance
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
Filtration
Signals: flow rate, pressure Prognosis Link
21 Kaggle - Production Plant Data for Condition Monitoring
von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018). Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems. Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRPCMS Production system
Productionline
Signals: anonymized Fault detection Link
22 Kaggle - Prognosis based on Varying Data Quality
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
Filtration
Signals: dust feed, flow rate, pressure Prognosis Link
23 Kaggle - Pump
Kaggle User: UnknownClass Process technology
Water Pump
Signals: anonymized Fault detection Link
24 Kaggle - Sensor Fault Detection
Schneider-Electric, France Electrical component
Temperature Sensor
Signals: temperature, voltage Fault detection Link
25 Kaggle - Versatile Production System
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
Versatile Production System
Signals: control parameters Fault detection Link
26 Mendeley - Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)
Lu, Jiahuan; Xiong, Rui; Tian, Jinpeng; Wang, Chenxu; Hsu, Chia-Wei; Tsou, Nien-Ti; Sun, Fengchun; Li, Ju (2021), “Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)”, Mendeley Data, V2, doi: 10.17632/kw34hhw7xg.2 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
27 Mendeley - Brushless DC motor
Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019). Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2 Drive technology
Brushless DC Motor
Signals: current, position, temperature Prognosis Link
28 Mendeley - Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells
Diao, Weiping (2019), “Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells”, Mendeley Data, V1,DOI: 10.17632/c35zbmn7j8.1 Electrical component
Battery
Signals: capacity Prognosis Link
29 Mendeley - Diesel Engine Faults
Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil Drive technology
Diesel Engine
Signals: pressure, temperature, vibration Diagnosis Link
30 Mendeley - HUST Bearing
Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3 Mechanical component
Bearing
Signals: vibration Diagnosis Link
31 Mendeley - Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell
Zuo, Jian (2024), “Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell”, Mendeley Data, V1, doi: 10.17632/w65jjt8v5w.1 Electrical component
Fuel Cell
Signals: unknown Prognosis Link
32 Mendeley - NMC cell 2600 mAh cyclic aging data
Burzyński, Damian; Kasprzyk, Leszek (2021), “NMC cell 2600 mAh cyclic aging data”, Mendeley Data, V1, doi: 10.17632/k6v83s2xdm.1 Electrical component
Battery
Signals: current, temperature Prognosis Link
33 Mendeley - Run-to-Failure Vibration Dataset of Self-Aligning Double-Row Ball Bearings
Gabrielli, Alberto; Battarra, Mattia; Mucchi, Emiliano; Dalpiaz, Giorgio (2024). Physics-based prognostics of rolling-element bearings: The equivalent damaged volume algorithm, Mechanical Systems and Signal Processing, Volume 215, 2024, 111435, doi: 10.1016/j.ymssp.2024.111435. Mechanical component
Bearing
Signals: vibration Prognosis Link
34 MFPT - Condition Based Maintenance Fault
Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems Mechanical component
Bearing
Signals: vibration Diagnosis Link
35 NASA - Accelerated Battery Life Testing
Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F. (2023). Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks. Proccedings of the Annual Conference of the PHM Society, 15(1). doi: 10.36001/phmconf.2023.v15i1.3463 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
36 NASA - Bearing Data Set
J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. ”Bearing Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Mechanical component
Bearing
Signals: vibration Prognosis Link
37 NASA - Capacitor Electrical Stress
J. Renwick, C. Kulkarni, and J Celaya ”Capacitor Electrical Stress Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Electrolytic Capacitors
Signals: impedance, voltage Prognosis Link
38 NASA - Capacitor Electrical Stress 2
J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel ”Capacitor Electrical Stress Data Set - 2”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Electrolytic Capacitor
Signals: capacity, resistance Prognosis Link
39 NASA - CFRP Composites Data Set
Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang ”CFRP Composites Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Materials
Carbon Fibre-Reinforced Polymer Composite
Signals: operating condition, strain, vibration, x-ray Prognosis Link
40 NASA - HIRF Battery
C. Kulkarni, E. Hogge, C. Quach and K. Goebel ”HIRF Battery Data Set”, NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognosticdata-repository), NASA Ames Research Center, Moffett Field, CA Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
41 NASA - IGBT
J. Celaya, Phil Wysocki, and K. Goebel ”IGBT Accelerated Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Transistor
Signals: , current, temperature, voltage Prognosis Link
42 NASA - Li-ion Battery Aging Datasets
B. Saha and K. Goebel. ”Battery Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
43 NASA - Milling Data set
A. Agogino and K. Goebel. BEST lab, UC Berkeley. ”Milling Data Set ”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Manufacturing process
Milling
Signals: acoustic emission, current, vibration Prognosis Link
44 NASA - MOSFET Thermal Overstress Aging
J. R. Celaya, A. Saxena, S. Saha, and K. Goebel ”MOSFET Thermal Overstress Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Transistor
Signals: current, temperature, voltage Prognosis Link
45 NASA - Randomized Battery Usage Data Set
B. Bole, C. Kulkarni, and M. Daigle ”Randomized Battery Usage Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
46 NASA - Small Satellite Power Simulation
C. Kulkarni and A. Guarneros ”Small Satellite Power Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Battery
Signals: current, temperature, voltage Diagnosis Link
47 NASA - Turbofan engine degradation simulation data set
A. Saxena and K. Goebel. ”Turbofan Engine Degradation Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
Aircraft Engine*
Signals: anonymized, operating condition Prognosis Link
48 NASA - Turbofan engine degradation simulation data set 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
Aircraft Engine*
Signals: anonymized, operating condition Prognosis Link
49 NIST - Robot Arm Position Accuracy
National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division Robotic
Industrial Robot
Signals: current, position, speed Diagnosis Link
50 OpenEI - Evaluation of Building Fault
Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019) Building
Buildings
Signals: pressure, speed, temperature Diagnosis Link
51 OpenEI - Gearbox Fault Diagnosis
Yogesh Pandy, National Renewable Energy Laboratory Mechanical component
Gear
Signals: vibration Fault detection Link
52 Oxford - Oxford Battery Degradation Dataset
Birkl, C. (2017). Oxford Battery Degradation Dataset 1. University of Oxford. DOI: 10.5287/bodleian:KO2kdmYGg
Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
53 PHM Data Challenge 2008 - Turbofan
NASA Prognostics Center of Excellence Drive technology
Aircraft Engine*
Signals: anonymized, operating condition Prognosis Link
54 PHM Data Challenge 2009 - Gearbox Fault Detection
PHM Society, Gearbox fault detection data set, 2010 Mechanical component
Gear System (Gears, Bearings, Shafts)
Signals: speed, vibration Diagnosis Link
55 PHM Data Challenge 2010 - CNC milling machine cutters
X. Li, B.S. Lim, J.H. Zhou, S. Huang, S.J. Phua, K.C. Shaw, and M.J. Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798 Manufacturing process
CNC Milling Machine Cutters
Signals: acoustic emission, force, vibration Prognosis Link
56 PHM Data Challenge 2011 - Anemometer Fault Detection
Creators Unknown Mechanical component
Anemometer
Signals: direction, speed, temperature Fault detection Link
57 PHM Data Challenge 2013 - Unknown
Creators Unknown Unknown
unknown
Signals: anonymized Health assessment Link
58 PHM Data Challenge 2014 - Unknown
Creators Unknown Unknown
unknown
Signals: anonymized Health assessment Link
59 PHM Data Challenge 2015 - Plant Fault Detection
Creators Unknown Production system
Industrial Plant Monitoring
Signals: anonymized Diagnosis Link
60 PHM Data Challenge 2016 - Semiconductor CMP
Crystec Technology Trading GmbH Manufacturing process
Chemicalmechanical Planarization System
Signals: flow rate, operating condition, pressure, speed Prognosis Link
61 PHM Data Challenge 2017 - Bogie Vehicle
Creators Unknown Mechanical component
Bogie
Signals: vibration Diagnosis Link
62 PHM Data Challenge 2018 - Ion Mill in Waver Manufactoring
Kai Goebel (at this time at NASA) Production system
Ion Mill Etching Tool
Signals: current, pressure, speed Prognosis Link
63 PHM Data Challenge 2019 - Fatigue Cracks
Creators Unknown Materials
Aluminum Structure with Dynamic Loading
Signals: vibration Prognosis Link
64 PHM Data Challenge 2020 Europe - Filtration System
Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412. Process technology
Filtration
Signals: flow rate, pressure Prognosis Link
65 PHM Data Challenge 2021 - Turbofan 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Runto-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2 ) Drive technology
Aircraft Engine*
Signals: anonymized, operating condition Prognosis Link
66 PHM Data Challenge 2021 Europe: SCARA-robot
Swiss Centre for Electronics and Microtechnology (CSEM) Robotic
Industrial Robot
Signals: current, duration, heat slope, humidity, position, pressure, speed, temperature Diagnosis Link
67 PHM Data Challenge 2022 - Rock Drills
Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden Mechatronic system
Rock Drills
Signals: pressure Diagnosis Link
68 PHM Data Challenge 2022 Europe - PCB Production line
Danilo Giordano and Martino Trevisan, (2022). ”PHME Data Challenge”. European conference of the prognostics and health management society. / BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en Production system
PCB Production Line
Signals: inspection data Diagnosis Link
69 PHM Data Challenge 2023 - Gearbox
PHM Society, Gearbox data set, 2023 Mechanical component
Gear
Signals: speed, vibration Diagnosis Link
70 PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
Japan Aerospace Exploration Agency (JAXA) Drive technology
Propulsion System*
Signals: pressure Health assessment Link
71 PHM Data Challenge 2024 - Helicopter Turbine Engines
Prognostics and Health Management (PHM) Society Drive technology
Helicopter Turbine Engine
Signals: power, speed, temperature, torque Fault detection Link
72 PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
FEMTO-ST Institute, Besan¸con, France Mechanical component
Bearing
Signals: temperature, vibration Prognosis Link
73 PHM IEEE Data Challenge 2014 - Fuel Cell
FCLAB Research Federation (FR CNRS 3539, France) Electrical component
Fuel Cell
Signals: current, current density, flow rate, hygrometry, pressure, temperature, voltage Prognosis Link
74 PHM IEEE Data Challenge 2023 - Planetary Gearbox
Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC) Mechanical component
Planetary Gear
Signals: vibration Diagnosis Link
75 PHM Society - Electromechanical Ball Screw Drive
GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects AgencyEnergy (ARPA-E), U.S. Department of Energy, under Award Number DEAR0001290 Mechatronic system
Electromechanical Ball Screw Drive*
Signals: current, position, speed, torque Diagnosis Link
76 ResearchGate - Driveline Unbalanced Shaft
Giacomo Barbieri and David SanchezLondono and Laura Cattaneo and Luca Fumagalli and David Romero Mechanical component
Unbalanced shaft (driveline)
Signals: vibration Diagnosis Link
77 SDOL - Diagnostics 101 bearing data
Kim, S., An, D., Choi, J-H. (2020). Diagnostics 101: A Tutorial for Fault Diagnostics of Rolling Element Bearing Using Envelope Analysis in MATLAB. Applied Sciences. 10(20):7302. doi:10.3390/app10207302 Mechanical component
Bearing
Signals: vibration Diagnosis Link
78 SDOL - HS Gear
Creators Unknown Mechanical component
Bearing
Signals: vibration Fault detection Link
79 SDOL - KAU Gear
Creators Unknown Mechanical component
Bearing
Signals: encoder Diagnosis Link
80 Toyota Research Institute - Battery Cycle Life
Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-0190356-8 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
81 UBFC - AMPERE Detection and diagnostics of rotor and stator faults in rotating machines
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06-03 Drive technology
Squirrel cage motor (Rotor, Stator, Bearing)
Signals: current, speed, vibration, voltage Diagnosis Link
82 UBFC - LASPI Detection and diagnostics of gearbox faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06 . Mechanical component
Bearing + Gear
Signals: current, vibration, voltage Diagnosis Link
83 UBFC - METALLICADOUR Detection and diagnostics of multi-axis robot faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire G´enie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère). Robotic
Multi-axes robot (cutting tool + axes drifts)
Signals: current, force, position, torque, vibration Diagnosis Link
84 UCI - AI4I 2020 Predictive Maintenance
Matzka, S. (2020). “Explainable Artificial Intelligence for Predictive Maintenance Applications,” Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp. 69-74, doi: 10.1109/AI4I49448.2020.00023 . Production system
Manufacturing Machine
Signals: speed, temperature, torque, wear Diagnosis Link
85 UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University Drive technology
Naval Propulsion Plant*
Signals: flow rate, injection control, pressure, speed, temperature, torque Prognosis Link
86 UCI - Condition monitoring of hydraulic systems Data Set
ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany Mechatronic system
Hydraulic System
Signals: efficiency, flow rate, power, pressure, temperature, vibration Diagnosis Link
87 UCI - Mechanical Analysis 1990
F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy Mechatronic system
Electromechanical Devices
Signals: speed Diagnosis Link
88 UCI - Robot Execution Failures Data Set
Universidade Nova de Lisboa, Monte da Caparica, Portugal Robotic
Industrial Robot
Signals: force, torque Diagnosis Link
89 UCI - Steel Plates Faults
Semeion Research Center of Sciences of Communication - Rome, Italy Materials
Steel Plates Faults
Signals: geometry, inspection data, operating condition Diagnosis Link
90 Uni Lulea - Vibration Signals Wind Turbines
Lulea University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements) Mechanical component
Bearing (Wind Turbines)
Signals: vibration Fault detection Link
91 Uni Paderborn KAt - Bearing Damage
University Paderborn Kat Mechanical component
Bearing
Signals: speed, temperature, torque Diagnosis Link
92 Virkler - Fatigue Crack Propagation
Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153. Materials
Aluminum Plate
Signals: crack length Prognosis Link
93 WASEDA - Fault Detection and Classification
De Bruijn, B., Nguyen, T. A., Bucur, D., & Tei, K. (2016). Benchmark datasets for fault detection and classification in sensor data. In A. Ahrens, O. Postolache, & C. Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Confererence on Sensor Networks, pp. 185-195. . SciTePress. https://doi.org/10.5220/0005637901850195 Electrical component
Sensors
Signals: light, temperature Diagnosis Link
94 Zenodo - Ball bearings subjected to time-varying load and speed conditions
Aimiyekagbon, O. K. (2024). Run-tofailure data set of ball bearings subjected to time-varying load and speed conditions. Zenodo, doi: 10.5281/zenodo.10805043, University Paderborn Mechanical component
Bearing
Signals: temperature, vibration Prognosis Link
95 Zenodo - Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation
Jiangong Zhu, Yixiu Wang, Yuan Huang, R. Bhushan Gopaluni, Yankai Cao, Michael Heere, Martin J. Mühlbauer, Liuda Mereacre, Haifeng Dai, Xinhua Liu, Anatoliy Senyshyn, Xuezhe Wei, Michael Knapp, & Helmut Ehrenberg. (2022). Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation [Data set]. Zenodo. https://doi.org/10.5281/zenodo.6405084 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
96 Zenodo - Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning
Zhang, Y., Tang, Q., Zhang, Y., Wang, J., Stimming, U., & Lee, A. A. (2020). Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning [Data set]. Zenodo. https://doi.org/10.5281/zenodo.3633835 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
97 Zenodo - Predictive maintenance dataset
Huawei Munich Research Center Mechatronic system
Elevator
Signals: humidity, vibration Prognosis Link
98 Zenodo - TBSI Sunwoda Battery Dataset
Shengyu, T. (2024). TBSI Sunwoda Battery Dataset [Data set]. Zenodo. https://doi.org/10.5281/zenodo.10715209 Electrical component
Battery
Signals: current, voltage Prognosis Link