### v3

# Fabian Mauthe

Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation [6]: Fabian Mauthe, Luca Steinmann, Moritz Neu, and Peter Zeiler. Overview and analysis of publicly available degradation data sets for tasks within prognostics and health management. In Proceedings of the 35th European Safety and Reliability Conference and the 33rd Society for Risk Analysis Europe Conference, pages 945–952, Singapore, 2025. Research Publishing. thanks: Citation [2]: Simon Hagmeyer, Fabian Mauthe, and Peter Zeiler. Creation of publicly available data sets for prognostics and diagnostics addressing data scenarios relevant to industrial applications. International Journal of Prognostics and Health Management, 12(2):1–20, 2021.
Fabian Mauthe
Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
Goeppingen, Germany
fabian.mauthe@hs-esslingen.de
   Christopher Braun
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Stuttgart, Germany
christopher.braun@iff.uni-stuttgart.de
   Julian Raible
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Stuttgart, Germany
julian.raible@iff.uni-stuttgart.de
   Peter Zeiler
Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
Goeppingen, Germany
peter.zeiler@hs-esslingen.de
   Marco F. Huber
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Fraunhofer Institute of Manufacturing Engineering and Automation IPA
Stuttgart, Germany
marco.huber@ieee.org
1 Prognostics and Health Management
In the realm of modern manufacturing, significant emphasis is placed on improving the reliability, performance, and service life of complex engineering systems, which has catalyzed the emergence of Prognostics and Health Management (PHM) as an essential discipline in the past two decades [10, 4]. PHM includes a range of methods dedicated to the early detection, prognosis, and mitigation of impending faults or failures in ESs, thereby facilitating sophisticated maintenance strategies and enhancing operational efficiency [1].
Central to the efficacy of PHM methods is the acquisition and analysis of degradation data, which encapsulates the evolving health condition of ESs over time [1]. Degradation data serves as a rich source of information, offering invaluable insights into the underlying degradation processes, failure modes, and performance trends of ESs [4, 2]. The availability of appropriate degradation data remains a significant challenge in both industrial practice and academic research [10]. Consequently, publicly available degradation data sets are of considerable value. They facilitate the development and empirical benchmarking of PHM methodologies and serve as a basis for demonstrating the practical applicability of the proposed approaches [5, 11, 7].
Although publicly available degradation data sets are of significant importance, the literature addresses this subject to a limited extent. Overviews often consider a restricted range of sources and platforms or only specific applications, resulting in constrained summaries [8, 9]. Analyses usually emphasize the amount of data or data quality, overlooking specific PHM aspects [3]. As a result, the search process for suitable data sets is often very time-consuming for users of PHM methods [6]. Therefore, the objective of this work is to provide a comprehensive overview of publicly available degradation data sets and a detailed PHM-specific analysis. To do this, the analysis presented in [6] is used. This PHM-specific analysis covers aspects that are pertinent to both users of PHM methods and of corresponding data sets. These aspects comprise the application represented in the data set, the originating domain of this application, the specific task the data set is designed to address, and the types of signals contained in the data set [6].
In the remainder of this work, the tasks of diagnostics and prognostics are listed in Section 2, and an overview of publicly available degradation data sets along with a PHM-specific analysis is presented in Section 3. Section 4 concludes the results of this work. The authors aim to update the overview as well as the analysis regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
2 Tasks of Diagnostics and Prognostics
Extensive and representative degradation data address a specific task of diagnostics and prognostics within PHM. Based on [10], [9], and [2], these tasks can be subdivided as follows:
  • Fault detection/anomaly detection: Detect a fault state or anomaly in an ES without considering its root cause. This results in a binary classification problem with the states fault or no fault.
  • Diagnosis: Assign one or more causes to a detected fault state.
  • Health assessment: Assess the state of degradation or the current risk of failure of an ES based on its current condition.
  • Prognosis: Prediction of future degradation behavior or the current Remaining Useful Life (RUL).
This list of tasks can also be interpreted as a sequential framework for implementing a comprehensive PHM application, starting with fault detection and ending with prognosis [10]. This can be illustrated by the following example of rolling bearing failure: Fault detection identifies an abnormal vibration pattern. Diagnosis then determines the root cause, for instance, a specific bearing component. The health assessment quantifies how this fault affects the bearing’s condition, while the prognosis predicts the bearing’s RUL. [6]
3 Data Sets for Diagnostics and Prognostics
3.1 Overview
Due to regular data challenges and several research activities of the PHM community in recent years, new data sets are continuously being published. Therefore, the overview by Mauthe et al. [6] (which is already based on the overview by Hagmeyer et al. [2]) is updated using a similar procedure, i.e., the identical scope of platforms and sources of degradation data sets is considered [2]. As a result, the overview is extended to 110 data sets in total. In accordance with [2] and [6], data sets that focus solely on process quality without including the degradation state are excluded from consideration. Table Appendix A: Overview of Publicly Available Data Sets in Appendix A: Overview of Publicly Available Data Sets contains the extended overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically.
3.2 Data Set Analysis
The detailed PHM-specific analysis of the data sets is based on the procedure presented in [6]. The analysis is intended to shorten the time-consuming search for suitable data sets for its users. In particular, the focus of this analysis is on the following PHM-specific aspects [6]:
  • Application: The application within a data set refers to the object of consideration from which the data originate. Thus, an application can represent a single component, such as a bearing, or an entire system, such as a production line.
  • Domain: The domain from which the application originates. The domains associated with the previous examples would then be, for example, mechanical components or production systems.
  • Task: The task, which is addressed by the corresponding data set. A distinction is made among the tasks described in Section 2: fault detection, diagnosis, health assessment, and prognosis.
  • Signal: The types of signals included in the data set. This refers to signals that offer insight into the health state of the specific application or can be derived from it and used within PHM methods.
Based on these PHM-specific aspects, Mauthe et al. [6] developed a taxonomy that is tailored to the classification of data sets in a PHM-specific manner. This taxonomy categorizes all data sets based on their application and domain of origin, the task addressed by the data set, and the relevant signals contained in the data set. The results using this taxonomy for the 110 publicly available data sets are provided in Table 1 in Appendix B: Classification of Publicly Available Data Sets and in Table 2 in Appendix C: Signals within Publicly Available Data Sets. The respective assignment of the PHM-specific aspects (application, domain, task, and signal) to the individual data sets can also be found in Table Appendix A: Overview of Publicly Available Data Sets in Appendix A: Overview of Publicly Available Data Sets.
Domains and applications:
Applying the presented taxonomy, the 110 publicly available degradation data sets are classified into 11 domains (see Table 1). The mechanical component domain contains the most data sets (26), followed by the electrical component domain (25) and the drive technology domain (12). Apart from these three largest domains, the remaining domains contain between two and eight data sets. Moreover, five data sets have unknown applications. As the respective data sets can still be used to apply PHM methods, they are taken into account and assigned to the unknown domain.
In total, 43 different applications are included in the 11 domains, covering a variety of degradation processes. However, as these applications are spread across 110 data sets, the number of available data sets per application is limited. This is reflected in a more in-depth consideration of Table 1: The applications bearing (17 data sets) and battery (16 data sets) are considered most frequently, originating from the two main domains, mechanical component and electrical component, respectively. The next most frequent applications are aircraft engine, filtration, and production line, each occurring in five data sets. The limited number of data sets per application is further highlighted by the mechatronic system domain, which contains the most different applications (eight), with only one data set per application. Overall, only nine applications are covered by three or more data sets.
Considered tasks within data sets:
The diagnostic and prognostic tasks of the data sets [^1] Certain data sets may be assigned to several tasks. As these tasks build on each other, the highest-ranking task is assigned to a respective data set. Therefore, for each data set, only one task is assigned., as introduced in Section 2, are shown in the third column of Table 1. With 54 assigned data sets, the prognosis task is the most represented, followed by diagnosis with 35 entries, whereas fault detection is assigned to 18 data sets. As in the previous classification, the transition from diagnosis to health assessment, as well as the transition to prognosis, is often not distinct. Accordingly, only three data sets explicitly address the task of health assessment.
For prognosis, the electrical component domain contributes 22 out of 54 data sets, of which 15 address battery applications. The mechanical component and drive technology domains follow with eight and seven prognosis data sets, respectively. Bearing is the second most common application for prognosis (seven data sets), while aircraft engine and filtration each occur in five prognosis data sets. In total, 20 distinct applications are available for prognosis.
For diagnosis, the mechanical component domain contains the most data sets (13 out of 35), while the robotic and mechatronic system domains contain five diagnosis data sets each, and the production system domain contains three. The bearing application is the most frequently used for diagnosis (eight data sets), followed by articulated robot applications (four) and both gear and production line applications (three each). In general, 21 different applications are considered in the 35 data sets for diagnosis, indicating a wide variety of applications.
For fault detection, the mechanical component and production system domains contribute the largest shares, with five and four data sets, respectively. Drive technology and manufacturing process contribute two fault detection data sets each, and the unknown domain contains two fault detection data sets without a specified application. Regarding applications, only bearing, milling, and production line occur in two data sets each; all remaining applications occur only once.
Signals within data sets:
Table 2 summarizes the 22 signals used in at least two of the given data sets. Note that Table 2 only lists signals that are measured directly. Values calculated from these, such as capacity or power, are excluded. Nevertheless, if only calculated signals are included in a data set, they are still listed. Also summarized are signals of a comparable nature, such as speed, including rotation and velocity.
Vibration appears most frequently, occurring in 34 data sets, followed by current and temperature with 33 data sets each, and voltage with 25 data sets. These signals are typically measured within electrical and mechanical component domains (e.g., battery and bearing applications) and therefore are most prevalent in the analyzed data sets. Speed and pressure are also common, appearing in 20 and 18 data sets, respectively. Anonymized signals and flow rate occur in ten data sets each; however, anonymized signals permit specific observations mainly within a data set or application, limiting their utility for general analysis. Regarding the availability of data sets per signal, a limitation similar to applications can be observed: ten of the 22 signals occur in four or fewer data sets.
Across tasks, signal occurrences show characteristic differences. For prognosis, current (22) and temperature (21) are most prevalent, followed by voltage (19) and vibration (14). For diagnosis, vibration dominates (16), followed by speed (11), current (nine), and temperature (eight), indicating that these signals are also relevant beyond purely electrical component data sets.
4 Conclusion
This work updates an overview of publicly available degradation data sets for tasks of diagnostics and prognostics within PHM and structures them using a dedicated taxonomy. The classification makes the current coverage across domains, applications, and tasks explicit and highlights where publicly available data are concentrated versus where only sparse coverage exists. Overall, the results underline that benchmarking is well supported for a few frequently studied use cases and signal modalities, while broader generalization and task-specific evaluations remain limited by uneven coverage. Maintaining and extending this overview will help track how coverage evolves and where additional publicly available data would have the greatest impact.
Appendix A: Overview of Publicly Available Data Sets
The designation of the data sets is divided into two parts, consisting of the source or platform and the title under which they were published. In addition, the overview contains the following columns:
  • Domain, Application, and Signals: The associated domain from which the data set originates and the respective application within this domain. In addition, the signals present in the respective data set are listed.
  • Task: The corresponding task within PHM addressable with this data set.
  • URL: The respective hyperlink to the source or platform hosting the data set.
Overview of publicly available data sets
No. Data Set Designation
Reference to Data Origin
Domain
Application and Signal
Task URL
[…]
No. Data Set Designation
Reference to Data Origin
Domain
Application and Signal
Task URL
[…]
Trad, Khiem (2021): Lifecycle ageing tests on commercial 18650 Li ion cell @ 25°C and 45°C. Version 1. 4TU.ResearchData. dataset. https://doi.org/10.4121/13739296.v1
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
2 4TU - Motor Current and Vibration Monitoring Dataset
Bruinsma, Sietze; Geertsma, Rinze; Loendersloot, Richard; Tinga, Tiedo (2024): Motor Current and Vibration Monitoring Dataset for various Faults in an E-motor-driven Centrifugal Pump. doi: 10.4121/2b61183ec14f-4131-829b-cc4822c369d0.v3
Process technology
Centrifugal pumps
Signals: current, vibration, voltage
Diagnosis Link
3 AIDAR Lab - Air Compressor
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2016). Intelligent Condition Based Monitoring using Acoustic Signals for Air Compressors, IEEE Transactions on Reliability, vol. 65, no. 1, pp. 291-309.
Mechatronic system
Air Compressor
Signals: acoustic emission
Diagnosis Link
4 AIDAR Lab - Drill Bit
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26.
Manufacturing process
Drill bit
Signals: vibration
Diagnosis Link
5 Aramis - Data Challenge ESREL2020PSAM15
Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France).
Unknown
Simulation (time-continuous stochastic process)*
Signals: vibration
Fault detection Link
6 Backblaze - Hard Drive Stats
Backblaze, Inc.
Drive technology
Disk Drives
Signals: temperature
Fault detection Link
7 Bearing Data Center - Fan and Bearing
Case Western Reserve University, USA
Mechanical component
Bearing
Signals: acoustic emission
Diagnosis Link
8 Calce - Battery Data Repository
CALCE Center for Advanced Life Cycle Engineering, University of Maryland, website: https://calce.umd.edu/battery-data
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
9 ETH Zurich - Run-to-Failure High-voltage Circuit Breaker Mechanical Test Dataset
Hsu, C.-C. (2024). Run-to-failure high-voltage circuit breaker mechanical test dataset. ETH Zürich Research Collection. DOI: 10.3929/ethz-b-000676480
Electrical component
Circuit breaker (high-voltage)
Signals: current, vibration
Prognosis Link
10 GitHub - Motor current milling machine
Apurv Rajeshkumar Darji (Developer), Dr. Mustafa Demetgül (Academic Supervisor)
Manufacturing process
Milling machine (motor current)
Signals: current
Fault detection Link
11 GitHub - Predictive Maintenance using PySpark
GitHub User: linya9191
Simulation
Simulation*
Signals: unknown
Fault detection Link
12 GitHub - XJTU-SY Bearing Datasets
Biao Wang, Yaguo Lei, Naipeng Li, Ningbo Li (2020). A Hybrid Prognostics Approach for Estimating Remaining Useful Life of Rolling Element Bearings, IEEE Transactions on Reliability, vol. 69, no. 1, pp. 401-412. DOI: 10.1109/TR.2018.2882682.
Mechanical component
Bearing
Signals: vibration
Prognosis Link
13 Harvard Dataverse - Wind Turbine Main Bearing Fatigue Life Prediction
Yucesan, Y. (2019). Wind Turbine Main Bearing Fatigue Life Prediction with PINN [Data set]. Harvard Dataverse. https://doi.org/10.7910/DVN/ENNXLZ
Mechanical component
Bearing
Signals: speed, temperature
Prognosis Link
14 Kaggle - Air pressure system failures in Scania trucks
Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm
Mechatronic system
Air Pressure System
Signals: anonymized
Fault detection Link
15 Kaggle - Bearings with Varying Degradation Behaviors
Mauthe, F.; Hagmeyer, S.; Zeiler, P. (2025). Holistic simulation model of the temporal degradation of rolling bearings. In E. B. Abrahamsen, T. Aven, F. Bouder, R. Flage, and M. Ylönen (Eds.), Proceedings of the 35th European Safety and Reliability Conference and the 33rd Society for Risk Analysis Europe Conference. Research Publishing. DOI: 10.3850/978-981-94-3281-3_ESREL-SRA-E2025-P8028-cd
Mechanical component
Bearing (simulated)
Signals: vibration
Prognosis Link
16 Kaggle - CNC Mill Tool Wear
System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan
Manufacturing process
Mill Tool
Signals: acceleration, current, feedrate, position, speed, voltage
Fault detection Link
17 Kaggle - Condition Data with Random Recording Time
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Signals: flow rate, pressure
Prognosis Link
18 Kaggle - E-coating ultrafiltration maintenance
Tinsley Bridge Limited, Sheffield, United Kingdom
Manufacturing process
Electrophoresis Painting
Signals: flow rate, pressure, temperature
Prognosis Link
19 Kaggle - Genesis Demonstrator
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Robotic
Pick-and-Place Demonstrator
Signals: acceleration, current, force, speed
Diagnosis Link
20 Kaggle - Microsoft Azure Predictive Maintenance
Azure AI Notebooks for Predictive Maintenance
Unknown
Unknown Machine
Signals: pressure, speed, vibration, voltage
Prognosis Link
21 Kaggle - One Year Industrial Component Degradation
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
Shrink-Wrapper
Signals: position, speed, torque
Fault detection Link
22 Kaggle - Predictive Maintenance 1
Creators Yuan Yao and Zhao Yuqi
Production system
Log Data
Signals: unknown
Fault detection Link
23 Kaggle - Preventive to Predictive Maintenance
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Signals: flow rate, pressure
Prognosis Link
24 Kaggle - Production Plant Data for Condition Monitoring
von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018). Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems. Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRPCMS
Production system
Production line
Signals: anonymized
Fault detection Link
25 Kaggle - Prognosis based on Varying Data Quality
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Signals: dust feed, flow rate, pressure
Prognosis Link
26 Kaggle - Pump
Kaggle User: UnknownClass
Process technology
Water Pump
Signals: anonymized
Fault detection Link
27 Kaggle - Sensor Fault Detection
Schneider-Electric, France
Electrical component
Temperature Sensor
Signals: temperature, voltage
Fault detection Link
28 Kaggle - Similar System Data Set for Condition Prognosis
Braig, M.; Zeiler, P. (2025): A Study on Using Transfer Learning to Utilize Information from Similar Systems for Data-Driven Condition Diagnosis and Prognosis. IEEE Access, Volume 13, pp. 98485-98503, DOI: 10.1109/ACCESS.2025.3576435
Process technology
Filtration
Signals: flow rate, pressure
Prognosis Link
29 Kaggle - Similar System Data Set for Fault Diagnosis
Braig, M.; Zeiler, P. (2025): A Study on Using Transfer Learning to Utilize Information from Similar Systems for Data-Driven Condition Diagnosis and Prognosis. IEEE Access, Volume 13, pp. 98485-98503, DOI: 10.1109/ACCESS.2025.3576435
Mechanical component
Bearing
Signals: vibration
Diagnosis Link
30 Kaggle - Versatile Production System
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
Versatile Production System
Signals: control parameters
Fault detection Link
31 Karlsruhe Institute of Technology - Ball Screw Drive Surface Defect Dataset
Schlagenhauf, T. (2021): Ball Screw Drive Surface Defect Dataset for Classification. Karlsruhe Institute of Technology (KIT). DOI: 10.5445/IR/1000133819
Mechanical component
Ball screw drive
Signals: image
Fault detection Link
32 Karlsruhe Institute of Technology - Industrial Machine Tool Element Surface Defect Dataset
Schlagenhauf, T.; Landwehr, M.; Fleischer, J. (2021): Industrial Machine Tool Element Surface Defect Dataset. Karlsruhe Institute of Technology (KIT). DOI: 10.5445/IR/1000129520
Mechanical component
Ball screw drive
Signals: image
Prognosis Link
33 Mendeley - Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)
Lu, Jiahuan; Xiong, Rui; Tian, Jinpeng; Wang, Chenxu; Hsu, Chia-Wei; Tsou, Nien-Ti; Sun, Fengchun; Li, Ju (2021), “Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)”, Mendeley Data, V2, doi: 10.17632/kw34hhw7xg.2
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
34 Mendeley - Brushless DC motor
Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019). Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58bdhy2df.2
Drive technology
Brushless DC Motor
Signals: current, position, temperature
Prognosis Link
35 Mendeley - Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells
Diao, Weiping (2019), “Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells”, Mendeley Data, V1,DOI: 10.17632/c35zbmn7j8.1
Electrical component
Battery
Signals: capacity
Prognosis Link
36 Mendeley - Diesel Engine Faults
Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil
Drive technology
Diesel Engine
Signals: pressure, temperature, vibration
Diagnosis Link
37 Mendeley - HUST Bearing
Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3
Mechanical component
Bearing
Signals: vibration
Diagnosis Link
38 Mendeley - Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell
Zuo, Jian (2024), “Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell”, Mendeley Data, V1, doi: 10.17632/w65jjt8v5w.1
Electrical component
Fuel Cell
Signals: unknown
Prognosis Link
39 Mendeley - NMC cell 2600 mAh cyclic aging data
Burzyński, Damian; Kasprzyk, Leszek (2021), “NMC cell 2600 mAh cyclic aging data”, Mendeley Data, V1, doi: 10.17632/k6v83s2xdm.1
Electrical component
Battery
Signals: current, temperature
Prognosis Link
40 Mendeley - Run-to-Failure Vibration Dataset of Self-Aligning Double-Row Ball Bearings
Gabrielli, Alberto; Battarra, Mattia; Mucchi, Emiliano; Dalpiaz, Giorgio (2024). Physics-based prognostics of rolling-element bearings: The equivalent damaged volume algorithm, Mechanical Systems and Signal Processing, Volume 215, 2024, 111435, doi: 10.1016/j.ymssp.2024.111435.
Mechanical component
Bearing
Signals: vibration
Prognosis Link
41 MFPT - Condition Based Maintenance Fault
Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems
Mechanical component
Bearing
Signals: vibration
Diagnosis Link
42 NASA - Accelerated Battery Life Testing
Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F. (2023). Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks. Proceedings of the Annual Conference of the PHM Society, 15(1). doi: 10.36001/phmconf.2023.v15i1.3463
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
43 NASA - Bearing Data Set
J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. ”Bearing Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Mechanical component
Bearing
Signals: vibration
Prognosis Link
44 NASA - Capacitor Electrical Stress
J. Renwick, C. Kulkarni, and J Celaya ”Capacitor Electrical Stress Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Electrolytic Capacitors
Signals: impedance, voltage
Prognosis Link
45 NASA - Capacitor Electrical Stress 2
J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel ”Capacitor Electrical Stress Data Set - 2”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Electrolytic Capacitor
Signals: capacity, resistance
Prognosis Link
46 NASA - CFRP Composites Data Set
Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang ”CFRP Composites Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Materials
Carbon Fibre-Reinforced Polymer Composite
Signals: operating condition, strain, vibration, x-ray
Prognosis Link
47 NASA - HIRF Battery
C. Kulkarni, E. Hogge, C. Quach and K. Goebel ”HIRF Battery Data Set”, NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognosticdata-repository), NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
48 NASA - IGBT
J. Celaya, Phil Wysocki, and K. Goebel ”IGBT Accelerated Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Transistor
Signals: current, temperature, voltage
Prognosis Link
49 NASA - Li-ion Battery Aging Datasets
B. Saha and K. Goebel. ”Battery Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
50 NASA - Milling Data set
A. Agogino and K. Goebel. BEST lab, UC Berkeley. ”Milling Data Set ”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Manufacturing process
Milling
Signals: acoustic emission, current, vibration
Prognosis Link
51 NASA - MOSFET Thermal Overstress Aging
J. R. Celaya, A. Saxena, S. Saha, and K. Goebel ”MOSFET Thermal Overstress Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Transistor
Signals: current, temperature, voltage
Prognosis Link
52 NASA - Randomized Battery Usage Data Set
B. Bole, C. Kulkarni, and M. Daigle ”Randomized Battery Usage Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
53 NASA - Small Satellite Power Simulation
C. Kulkarni and A. Guarneros ”Small Satellite Power Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Signals: current, temperature, voltage
Diagnosis Link
54 NASA - Turbofan engine degradation simulation data set
A. Saxena and K. Goebel. ”Turbofan Engine Degradation Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
Aircraft Engine*
Signals: anonymized, operating condition
Prognosis Link
55 NASA - Turbofan engine degradation simulation data set 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
Aircraft Engine*
Signals: anonymized, operating condition
Prognosis Link
56 NIST - Robot Arm Position Accuracy
National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division
Robotic
Industrial Robot
Signals: current, position, speed
Diagnosis Link
57 OpenEI - Evaluation of Building Fault
Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019)
Building
Buildings
Signals: pressure, speed, temperature
Diagnosis Link
58 OpenEI - Gearbox Fault Diagnosis
Yogesh Pandy, National Renewable Energy Laboratory
Mechanical component
Gear
Signals: vibration
Fault detection Link
59 OSF - Second-life lithium-ion battery aging dataset based on grid storage cycling
Khan, M. A., & Onori, S. (2025). Second-life lithium-ion battery aging dataset based on grid storage cycling. https://doi.org/10.17605/OSF.IO/8JNR5
Electrical component
Battery
Signals: current, impedance, voltage
Prognosis Link
60 Oxford - Oxford Battery Degradation Dataset
Birkl, C. (2017). Oxford Battery Degradation Dataset 1. University of Oxford. DOI: 10.5287/bodleian:KO2kdmYGg
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
61 PHM Data Challenge 2008 - Turbofan
NASA Prognostics Center of Excellence
Drive technology
Aircraft Engine*
Signals: anonymized, operating condition
Prognosis Link
62 PHM Data Challenge 2009 - Gearbox Fault Detection
PHM Society, Gearbox fault detection data set, 2010
Mechanical component
Gear System (Gears, Bearings, Shafts)
Signals: speed, vibration
Diagnosis Link
63 PHM Data Challenge 2010 - CNC milling machine cutters
X. Li, B.S. Lim, J.H. Zhou, S. Huang, S.J. Phua, K.C. Shaw, and M.J. Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798
Manufacturing process
CNC Milling Machine Cutters
Signals: acoustic emission, force, vibration
Prognosis Link
64 PHM Data Challenge 2011 - Anemometer Fault Detection
Creators Unknown
Mechanical component
Anemometer
Signals: direction, speed, temperature
Fault detection Link
65 PHM Data Challenge 2013 - Unknown
Creators Unknown
Unknown
unknown
Signals: anonymized
Health assessment Link
66 PHM Data Challenge 2014 - Unknown
Creators Unknown
Unknown
unknown
Signals: anonymized
Health assessment Link
67 PHM Data Challenge 2015 - Plant Fault Detection
Creators Unknown
Production system
Industrial Plant Monitoring
Signals: anonymized
Diagnosis Link
68 PHM Data Challenge 2016 - Semiconductor CMP
Crystec Technology Trading GmbH
Manufacturing process
Chemical mechanical planarization system
Signals: flow rate, operating condition, pressure, speed
Prognosis Link
69 PHM Data Challenge 2017 - Bogie Vehicle
Creators Unknown
Mechanical component
Bogie
Signals: vibration
Diagnosis Link
70 PHM Data Challenge 2018 - Ion Mill in Wafer Manufacturing
Kai Goebel (at this time at NASA)
Production system
Ion Mill Etching Tool
Signals: current, pressure, speed
Prognosis Link
71 PHM Data Challenge 2019 - Fatigue Cracks
He J, Guan X, Peng T, Liu Y, Saxena A, Celaya J, Goebel K (2013). A multi-feature integration method for fatigue crack detection and crack length estimation in riveted lap joints using Lamb waves. Smart Materials and Structures. 22(10):105007.
Materials
Aluminum Structure with Dynamic Loading
Signals: vibration
Prognosis Link
72 PHM Data Challenge 2020 Europe - Filtration System
Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412.
Process technology
Filtration
Signals: flow rate, pressure
Prognosis Link
73 PHM Data Challenge 2021 - Turbofan 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2 )
Drive technology
Aircraft Engine*
Signals: anonymized, operating condition
Prognosis Link
74 PHM Data Challenge 2021 Europe: SCARA-robot
Swiss Centre for Electronics and Microtechnology (CSEM)
Robotic
Industrial Robot
Signals: current, duration, heat slope, humidity, position, pressure, speed, temperature
Diagnosis Link
75 PHM Data Challenge 2022 - Rock Drills
Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden
Mechatronic system
Rock Drills
Signals: pressure
Diagnosis Link
76 PHM Data Challenge 2022 Europe - PCB Production line
Danilo Giordano and Martino Trevisan, (2022). ”PHME Data Challenge”. European conference of the prognostics and health management society. / BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en
Production system
PCB Production Line
Signals: inspection data
Diagnosis Link
77 PHM Data Challenge 2023 - Gearbox
PHM Society, Gearbox data set, 2023
Mechanical component
Gear
Signals: speed, vibration
Diagnosis Link
78 PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
Japan Aerospace Exploration Agency (JAXA)
Drive technology
Propulsion System*
Signals: pressure
Health assessment Link
79 PHM Data Challenge 2024 - Helicopter Turbine Engines
Prognostics and Health Management (PHM) Society
Drive technology
Helicopter Turbine Engine
Signals: power, speed, temperature, torque
Fault detection Link
80 PHM Data Challenge 2025 - Aircraft Engines
Creators Unknown, data generation: NASA AGTF30 Simulation – MATLAB Executable Steady-State Solver and Linearization Tool for the AGTF30 Engine Simulation (MEXLIN-AGTF30), https://software.nasa.gov/software/LEW-20688-1. See also https://software.nasa.gov/software/LEW-19717-1. MATLAB T-MATS Toolbox – https://ntrs.nasa.gov/api/citations/20180002976/downloads/20180002976.pdf
Drive technology
Aircraft Engine*
Signals: operation conditions, pressure
Prognosis Link
81 PHM Data Challenge 2025 Asia Pacific - Predicting Cutter Flank Wear
PHM Asia Pacific
Manufacturing process
Turning machine cutter flank wear
Signals: acoustic emission, vibration
Prognosis Link
82 PHM Data Challenge 2026 Europe - Subway ticket validation door
Soualhi, M. (2026): PHME Data Challenge. European conference of the prognostics and health management society
Mechatronic system
Validation door
Signals: current, position, voltage
Prognosis Link
83 PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
FEMTO-ST Institute, Besançon, France
Mechanical component
Bearing
Signals: temperature, vibration
Prognosis Link
84 PHM IEEE Data Challenge 2014 - Fuel Cell
FCLAB Research Federation (FR CNRS 3539, France)
Electrical component
Fuel Cell
Signals: current, current density, flow rate, hygrometry, pressure, temperature, voltage
Prognosis Link
85 PHM IEEE Data Challenge 2023 - Planetary Gearbox
Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC)
Mechanical component
Planetary Gear
Signals: vibration
Diagnosis Link
86 PHM Society - Electromechanical Ball Screw Drive
GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy, under Award Number DEAR0001290
Mechatronic system
Ball Screw Drive* (Electromechanical)
Signals: current, position, speed, torque
Diagnosis Link
87 ResearchGate - Driveline Unbalanced Shaft
Giacomo Barbieri and David Sanchez-Londono and Laura Cattaneo and Luca Fumagalli and David Romero
Mechanical component
Unbalanced shaft (driveline)
Signals: vibration
Diagnosis Link
88 SDOL - Diagnostics 101 bearing data
Kim, S., An, D., Choi, J-H. (2020). Diagnostics 101: A Tutorial for Fault Diagnostics of Rolling Element Bearing Using Envelope Analysis in MATLAB. Applied Sciences. 10(20):7302. doi:10.3390/app10207302
Mechanical component
Bearing
Signals: vibration
Diagnosis Link
89 SDOL - HS Gear
Creators Unknown
Mechanical component
Bearing
Signals: vibration
Fault detection Link
90 SDOL - KAU Gear
Creators Unknown
Mechanical component
Bearing
Signals: encoder
Diagnosis Link
91 Toyota Research Institute - Battery Cycle Life
Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-0190356-8
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
92 UBFC - AMPERE Detection and diagnostics of rotor and stator faults in rotating machines
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06-03
Drive technology
Squirrel cage motor (Rotor, Stator, Bearing)
Signals: current, speed, vibration, voltage
Diagnosis Link
93 UBFC - LASPI Detection and diagnostics of gearbox faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06 .
Mechanical component
Bearing + Gear
Signals: current, vibration, voltage
Diagnosis Link
94 UBFC - METALLICADOUR Detection and diagnostics of multi-axis robot faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère).
Robotic
Multi-axes robot (cutting tool + axes drifts)
Signals: current, force, position, torque, vibration
Diagnosis Link
95 UCI - AI4I 2020 Predictive Maintenance
Matzka, S. (2020). “Explainable Artificial Intelligence for Predictive Maintenance Applications,” Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp. 69-74, doi: 10.1109/AI4I49448.2020.00023 .
Production system
Manufacturing Machine
Signals: speed, temperature, torque, wear
Diagnosis Link
96 UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University
Drive technology
Naval Propulsion Plant*
Signals: flow rate, injection control, pressure, speed, temperature, torque
Prognosis Link
97 UCI - Condition monitoring of hydraulic systems Data Set
ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany
Mechatronic system
Hydraulic System
Signals: efficiency, flow rate, power, pressure, temperature, vibration
Diagnosis Link
98 UCI - Mechanical Analysis 1990
F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy
Mechatronic system
Electromechanical Devices
Signals: speed
Diagnosis Link
99 UCI - Robot Execution Failures Data Set
Universidade Nova de Lisboa, Monte da Caparica, Portugal
Robotic
Industrial Robot
Signals: force, torque
Diagnosis Link
100 UCI - Steel Plates Faults
Semeion Research Center of Sciences of Communication - Rome, Italy
Materials
Steel Plates Faults
Signals: geometry, inspection data, operating condition
Diagnosis Link
101 Uni Lulea - Vibration Signals Wind Turbines
Lulea University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements)
Mechanical component
Bearing (Wind Turbines)
Signals: vibration
Fault detection Link
102 Uni Paderborn KAt - Bearing Damage
University Paderborn Kat
Mechanical component
Bearing
Signals: speed, temperature, torque
Diagnosis Link
103 Virkler - Fatigue Crack Propagation
Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153.
Materials
Aluminum Plate
Signals: crack length
Prognosis Link
104 WASEDA - Fault Detection and Classification
De Bruijn, B., Nguyen, T. A., Bucur, D., & Tei, K. (2016). Benchmark datasets for fault detection and classification in sensor data. In A. Ahrens, O. Postolache, & C. Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Conference on Sensor Networks, pp. 185-195. SciTePress. https://doi.org/10.5220/0005637901850195
Electrical component
Sensors
Signals: light, temperature
Diagnosis Link
105 Zenodo - Ball bearings subjected to time-varying load and speed conditions
Aimiyekagbon, O. K. (2024). Run-to-failure data set of ball bearings subjected to time-varying load and speed conditions. Zenodo, doi: 10.5281/zenodo.10805043, University Paderborn
Mechanical component
Bearing
Signals: temperature, vibration
Prognosis Link
106 Zenodo - Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation
Jiangong Zhu, Yixiu Wang, Yuan Huang, R. Bhushan Gopaluni, Yankai Cao, Michael Heere, Martin J. Mühlbauer, Liuda Mereacre, Haifeng Dai, Xinhua Liu, Anatoliy Senyshyn, Xuezhe Wei, Michael Knapp, & Helmut Ehrenberg. (2022). Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation [Data set]. Zenodo. https://doi.org/10.5281/zenodo.6405084
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
107 Zenodo - Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning
Zhang, Y., Tang, Q., Zhang, Y., Wang, J., Stimming, U., & Lee, A. A. (2020). Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning [Data set]. Zenodo. https://doi.org/10.5281/zenodo.3633835
Electrical component
Battery
Signals: current, temperature, voltage
Prognosis Link
108 Zenodo - Monitoring data railway bridge Leuven
Kristof Maes, & Geert Lombaert. (2020). Monitoring data for railway bridge KW51 in Leuven, Belgium, before, during, and after retrofitting [Data set]. Zenodo. https://doi.org/10.5281/zenodo.3745914
Building
Railway bridge
Signals: acceleration, displacement, strain
Diagnosis Link
109 Zenodo - Predictive maintenance dataset
Huawei Munich Research Center
Mechatronic system
Elevator
Signals: humidity, vibration
Prognosis Link
110 Zenodo - TBSI Sunwoda Battery Dataset
Shengyu, T. (2024). TBSI Sunwoda Battery Dataset [Data set]. Zenodo. https://doi.org/10.5281/zenodo.10715209
Electrical component
Battery
Signals: current, voltage
Prognosis Link
Appendix B: Classification of Publicly Available Data Sets
[…]
Domain Application FD D HA P Sum across domain
Building Building 1 2
Bridge 1
Drive technology Aircraft engine* 5 12
Electric motor 1 1 1
Diesel engine 1
Helicopter engine 1
Propulsion system* 1 1
Electrical component Battery 1 15 25
Capacitor 2
Circuit breaker 1
Fuel cell 2
Sensor 1 1
Transistor 2
Manufacturing process Planarization system 1 8
Drilling 1
Electrophoresis painting 1
Milling 2 2
Turning 1
Material Aluminum plate 2 4
Polymer composite 1
Steel plate 1
Mechanical component Anemometer 1 26
Ball screw drive 1 1
[…]
Bogie 1
Gear 1 3
Shaft 1
Mechatronic system Air compressor 1 8
Air pressure system 1
Electromechanical ball screw* 1
Electromechanical device 1
Elevator 1
Hydraulic system 1
Rock drill 1
Validation door 1
Process technology Filtration 5 7
Pump 1 1
Production system Ion mill etching tool 1 8
Log data 1
Production line 2 3
Shrink-wrapper 1
Robotic Articulated robot 4 5
Linear robot 1
Unknown 2 2 1 5
all 18 35 3 54 110
Appendix C: Signals within Publicly Available Data Sets
Table 2: Signals used in at least two of the given data sets. Sorted by the sum of signal occurrences across the tasks of fault detection (FD), diagnosis (D), health assessment (HA), and prognosis (P).
Signal FD D HA P Sum across tasks
Vibration 4 16 0 14 34
Current 2 9 0 22 33
Temperature 4 8 0 21 33
Voltage 2 4 0 19 25
Speed 4 11 0 5 20
Pressure 0 5 1 12 18
Anonymized(1) 3 1 2 4 10
Flow Rate 0 1 0 9 10
Position 2 4 0 2 8
Torque 2 5 0 1 8
Operating Condition(2) 0 1 0 6 7
Acoustic Emission 0 2 0 3 5
Force 0 3 0 1 4
Acceleration(3) 1 2 0 0 3
Unknown(4) 2 0 0 1 3
Capacity 0 0 0 2 2
Humidity 0 1 0 1 2
Image 1 0 0 1 2
Impedance 0 0 0 2 2
Inspection Data 0 2 0 0 2
Power 1 1 0 0 2
Strain 0 1 0 1 2
(1): signal values changed/transformed,  (2): varying operating conditions during the life
(3): in terms of movement, not vibration, (4): signal (or sensor) type unknown